OpenStudy (anonymous):

Find the surface area of a cone with diameter 18 inches and slant height 22 inches. A. 452.47 in2 B. 703.04 in2 C. 650.31 in2 D. 876.5 in2

OpenStudy (anonymous):

plz help

OpenStudy (mathstudent55):

\( SA_cone = \pi rs + \pi r^2 \) r = radius = diameer/2 s = slant height

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it b?